OpenStudy (anonymous):

Prove the identity. cos^2x-sin^2x-1=1-2sin^2x

OpenStudy (anonymous):

something is wrong here..

OpenStudy (anonymous):

are you sure that this is the question?

OpenStudy (anonymous):

yes but we were thought to use theta in the place of x. there is not sign for it on my computer

OpenStudy (anonymous):

no this is not the problem here.. there is something wrong in the question it isnt right lol

OpenStudy (anonymous):

check if you copied it right

OpenStudy (anonymous):

oh lol this is what the problem shows here. cos squared x-sin squared x=1-2sin squared x.

OpenStudy (anonymous):

i wrote it out to see it that helps but that is how it is written.

OpenStudy (anonymous):

oh now its fine.. you see you wrote in the first place an extra -1 on the left side so: using cos^2x + sin^2x = 1 we get cos^2x = 1-sin^2x plug it into your equation 1-sin^2x -sin^2x = 1 -2sin^2x 1-2sin^2x = 1-2sin^2x done

OpenStudy (anonymous):

thank you